# PoLi-RL: A Point-to-List Reinforcement Learning Framework for Conditional Semantic Textual Similarity

### v1

@@ Abstract
Conditional Semantic Textual Similarity (C-STS) measures the semantic proximity between text segments under a specific condition, thereby overcoming the ambiguity inherent in traditional STS. However, existing methods are largely confined to discriminative models, failing to fully integrate recent breakthroughs in the NLP community concerning Large Language Models (LLMs) and Reinforcement Learning (RL). RL is a particularly well-suited paradigm for this task, as it can directly optimize the non-differentiable Spearman ranking metric and guide the reasoning process required by C-STS. However, we find that naively applying listwise RL fails to produce meaningful improvements, as the model is overwhelmed by complex, coarse-grained reward signals. To address this challenge, we introduce PoLi-RL, a novel Point-to-List Reinforcement Learning framework. PoLi-RL employs a two-stage curriculum: it first trains the model with simple pointwise rewards to establish fundamental scoring capabilities, then transitions to a hybrid reward that combines pointwise, pairwise, and listwise objectives to refine the model’s ability to discern subtle semantic distinctions. Crucially, we propose an innovative Parallel Slice Ranking Reward (PSRR) mechanism that computes ranking rewards in parallel slices, where each slice comprises same-indexed completions from different samples. This provides a precise, differentiated learning signal for each individual completion, enabling granular credit assignment and effective optimization. On the official C-STS benchmark, PoLi-RL achieves a Spearman correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture. As the first work to successfully apply RL to C-STS, our study introduces a powerful and precise paradigm for training LLMs on complex, ranking-based conditional judgment tasks.
11footnotetext: Equal Contribution. Work done during the internship at Tencent Youtu Lab.
@@ 1 Introduction
As a core research area in Computational Linguistics, Semantic Textual Similarity (STS) (Agirre et al., 2013) finds extensive applications across diverse scenarios, including topic modeling, dialogue systems, text summarization, and agent memory (Tang et al., 2025). However, traditional STS tasks exhibit inherent ambiguity because similarity definitions are often susceptible to observer bias. To address this limitation, the Conditional Semantic Textual Similarity (C-STS) task was developed (Deshpande et al., 2023). By incorporating an explicit natural language condition, C-STS enables more precise and objective similarity judgments, yet simultaneously imposes higher demands on a model’s reasoning capabilities. For instance, consider the following two text fragments: “A player is shooting from beyond the three-point line” and “A player is taking a free throw”. Under the condition “The activity of the player”, their similarity is high. However, under the condition “The player’s distance from the basket”, their similarity is low.
@@
Research on this nascent task has yielded three primary paradigms: Bi-encoder (Liu et al., 2025), Tri-encoder (Lin et al., 2024), and Cross-encoder (Li et al., 2024). The Cross-encoder architecture, which processes the text pair and the guiding condition simultaneously, is the most compatible with modern generative pre-trained models. Despite this, the integration of C-STS with LLMs remains in its early stages. Current LLM applications are limited to two main approaches: direct inference via few-shot prompting, where even state-of-the-art models struggle to achieve satisfactory results (Deshpande et al., 2023); and their use as feature extractors for generating text embeddings (Yamada and Zhang, 2025), which is an extension of the discriminative paradigm. To the best of our knowledge, no prior work has applied an end-to-end LLM-based cross-encoder to the C-STS task, nor has any integrated it with advanced training techniques like reinforcement learning (RL), leaving a significant research gap.
@@
This paper aims to fill this gap. We posit that incorporating RL into an LLM-based cross-encoder paradigm is a natural fit. This is reflected in two aspects: First, C-STS requires nuanced, scenario-based reasoning. For example, in the basketball case described earlier, to correctly assess similarity under the “distance” condition, the model must move beyond surface-level semantics to identify the underlying spatial relationship between ‘beyond the three-point line’ and ‘at the free throw line’, a process demanding strong abstraction and inference. RL, through its explicit reward signals, can more effectively guide the reasoning process of LLMs (Guo et al., 2025). Second, from an optimization standpoint, RL aligns closely with the task’s evaluation criteria. The Spearman correlation coefficient (Zar, 2005), a core evaluation metric of C-STS, is a non-differentiable measure of ranking quality. Traditional SFT methods can only indirectly and approximately optimize this objective through loss functions like Mean Squared Error (MSE) (Zhang and Li, 2024). In contrast, RL allows for the direct optimization of ranking-based reward functions that are designed to correlate strongly with the final Spearman metric, maintaining a higher degree of consistency with the final evaluation.
@@
However, a naive application of RL to this task presents significant challenges. As illustrated in Figure 1, our preliminary experiments indicate that directly applying a single listwise ranking reward (e.g., Spearman’s correlation coefficient) across an entire batch of completions does not show any advantages compared to the few-shot method. This approach suffers from two fundamental problems. First, the ranking objective is too complex for a model that has not yet learned the task’s fundamental scoring rules, often causing the training to stagnation. Second, a single reward computed across the entire batch is too coarse to provide precise credit assignment, as a few poor completions can unfairly penalize other good ones.
@@
Figure 1: Performance comparison of different strategies on the C-STS task. Directly applying listwise ranking rewards for RL does not significantly outperform the few-shot baseline. In contrast, both stages of our method (PoLi-RL) achieve substantial improvements, validating its effectiveness.
@@
To address these challenges, we propose PoLi-RL, a two-stage Point-to-List Reinforcement Learning framework. PoLi-RL features a two-stage curriculum to manage the complexity of the learning task. In the first stage, we use simple pointwise rewards to ground the model in the basic scoring rules of the task. Building on this foundation, the second stage introduces a richer, hybrid reward signal that combines a stable pointwise anchor with more nuanced pairwise and listwise ranking rewards, enabeling the model to discern subtle semantic differences.
@@
Furthermore, to resolve the problem of coarse-grained reward signals that arises from ranking all completions in a batch together, we innovatively introduce a Parallel Slice Ranking Reward (PSRR) mechanism, which utilizes a two-level decomposition. First, for a batch of input samples, the model generates $G$ completions for each. We then form $G$ “parallel slices”, where the $i$-th slice is composed of the $i$-th completion from every sample. Second, and more importantly, within each slice, rather than assigning a single reward, we compute the rank difference for each individual completion against its ideal rank. This two-level decomposition allows each of the $N\times G$ completions to receive a unique and precise reward that reflects its quality, thereby facilitating granular credit assignment and more effective optimization.
@@
  • To the best of our knowledge, this is the first work to propose an end-to-end, LLM-based cross-encoder for the C-STS task and the first to employ reinforcement learning for training in this domain.
  • We design and implement PoLi-RL [^1] The code is available at: https://github.com/ZBWpro/PoLi-RL., a novel two-stage training curriculum that addresses the complexity of C-STS by progressing from a simple pointwise reward to a more complex hybrid reward.
  • We propose the Parallel Slice Ranking Reward (PSRR) mechanism, which delivers precise and differentiated learning signals by computing ranking rewards within independent ‘parallel slices’ to enable granular credit assignment. This mechanism offers a generalizable strategy for other ranking and retrieval tasks involving multiple generation candidates.
  • On the official C-STS benchmark, PoLi-RL achieves a Spearman’s correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture and surpassing strong closed-source models, including GPT-4 (43.6) (Achiam et al., 2023). Our qualitative analysis further reveals our method’s advantages in understanding complex conditions.
@@ 2 Methodology
This section details our proposed strategy. We begin by formulating the C-STS task within an end-to-end, LLM-based cross-encoder paradigm in subsection 2.1. Then, in subsection 2.2, we map this task onto the mathematical framework of Reinforcement Learning and specifiy its optimization objective. Finally, in subsection 2.3, we provide a comprehensive description of our PoLi-RL framework, including its two-stage design and the innovative PSRR mechanism.
@@ 2.1 Problem Formulation
The core objective of C-STS is to learn a scoring function that accurately reflects the semantic similarity between two text segments under a specific condition. Formally, each C-STS data sample is defined as a tuple $x=(t_{1},t_{2},c,y)$, where $t_{1},t_{2}$ are two text segments, $c$ is the natural language condition, and $y\in[1,5]$ is the human-annotated similarity judgement on the Likert scale (Likert, 1932). Notably, the label $y$ corresponds to a fine-grained set of semantic criteria. According to the C-STS annotation guidelines, the meanings of the scores are as follows: (1) Completely dissimilar; (2) Thematically related but dissimilar; (3) Roughly equivalent, but with some important information differences; (4) Mostly equivalent, with some unimportant details differing; (5) Completely equivalent. This level of granularity demands that the model perform fine-grained reasoning beyond surface-level semantics, posing a significant challenge to its capabilities.
A unique characteristic of the C-STS dataset is its paired structure: samples are organized in adjacent pairs that share the same text segments $(t_{1},t_{2})$ but feature different conditions and maintain a deterministic ordinal relationship between their labels, i.e., $y_{\text{high}}\geq y_{\text{low}}$. This structure provides a solid foundation for our pairwise reward design, as detailed in subsection 2.3.
Our task is to train a scoring model $\pi_{\theta}$, parameterized by $\theta$. For each sample $x$, the model takes a unified prompt $p=\text{INS}(x)$ as input, which is constructed from the text segments, condition, and task-specific instructions (detailed prompt templates are provided in A.2), to generate an output sequence $o=\pi_{\theta}(p)$. From this sequence, we parse the final predicted score, $\tilde{y}=\text{Parse}(o)$. The overall training objective is to maximize the ranking consistency between the set of predicted scores $\{\tilde{y}_{i}\}_{i=1}^{N}$ and the ground-truth scores $\{y_{i}\}_{i=1}^{N}$ by optimizing the policy $\pi_{\theta}$, a process primarily measured by Spearman’s correlation coefficient. Since this metric is rank-based and non-differentiable, RL emerges as a more promising optimization paradigm than traditional supervised fine-tuning.
@@ 2.2 Reinforcement Learning for C-STS
We formulate the C-STS task as a Markov Decision Process (MDP), defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where the agent is the LLM policy $\pi_{\theta}$. The generation process is modeled as a sequence of decisions, where each step involves generating a single token. A state $s_{t}\in\mathcal{S}$ at timestep $t$ is the sequence of tokens generated so far, conditioned on the initial prompt, i.e., $s_{t}=(p,o_{<t})$. An action $a_{t}\in\mathcal{A}$ is the selection of the next token $o_{t}$ from the model’s vocabulary, governed by the policy $\pi_{\theta}(a_{t}|s_{t})$, which provides a probability distribution over all possible tokens. The transition function $\mathcal{T}$ is deterministic given a selected action: the next state $s_{t+1}$ is formed by appending the chosen token $a_{t}$ to the current state $s_{t}$. The stochasticity in the generation process arises from the sampling procedure, where actions are drawn from the policy distribution $\pi_{\theta}(a_{t}|s_{t})$ with configurable parameters such as temperature and top-p. We employ a terminal reward setting, where a reward $\mathcal{R}_{T}=\mathcal{R}(x,o)$ is given only after the entire sequence $o$ has been generated. Finally, the discount factor $\gamma$ is set to 1 to ensure that the terminal reward is backpropagated without decay to all actions that contributed to the final output. Based on this framework, the objective is to find the optimal parameters $\theta^{*}$ that maximize the expected reward over the data distribution $\mathcal{D}$:
$\theta^{*}=\underset{\theta}{\operatorname{arg\,max}}\mathbb{E}_{x\sim\mathcal{D},o\sim\pi_{\theta}(p)}[R(x,o)]$ (1)
To optimize this objective, we employ Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), an extension of GRPO (Shao et al., 2024) that introduces several key techniques for effective RL. For each sample $x$, the policy generates a set of $G$ completions $\{o_{i}\}_{i=1}^{G}$. A scalar reward $r_{i}=R(x,o_{i})$ is computed for each completion. The advantage $\hat{A}_{i}$ for each completion is then calculated by normalizing its reward against the statistics of the entire group’s rewards via Z-score normalization:
@@
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|p)}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left(\frac{\pi_{\theta}(o_{i,t}|p,o_{i,<t})}{[\pi_{\theta}(o_{i,t}|p,o_{i,<t})]_{\text{nograd}}}\hat{A}_{i,t}\right)\right]$ (3)
where $\cdot]_{\text{nograd}}$ denotes stopping the gradient flow through the denominator, ensuring that only the numerator is updated during backpropagation.

### v2

@@ Abstract
Conditional Semantic Textual Similarity (C-STS) measures the semantic proximity between text segments under a specific condition, thereby overcoming the ambiguity inherent in traditional STS. However, existing methods are largely confined to discriminative models, failing to fully leverage recent breakthroughs in the NLP community involving Large Language Models (LLMs) and Reinforcement Learning (RL). RL is a particularly well-suited paradigm for this task, as it can directly optimize the non-differentiable Spearman ranking metric and guide the reasoning process required by C-STS. Nevertheless, we find that naively applying listwise RL fails to produce meaningful improvements, as the model struggles with complex, coarse-grained reward signals, leading to optimization difficulties. To address this challenge, we introduce PoLi-RL, a novel Point-to-List Reinforcement Learning framework. PoLi-RL employs a two-stage curriculum: it first trains the model with a simple pointwise reward to establish fundamental scoring capabilities, then transitions to a hybrid reward that combines pointwise, pairwise, and listwise objectives to refine the model’s ability to discern subtle semantic distinctions. Crucially, we propose an innovative Parallel Slice Ranking Reward (PSRR) mechanism that computes ranking rewards in parallel slices, where each slice consists of completions with the same index from different samples. This provides a precise, differentiated learning signal for each individual completion, enabling granular credit assignment and effective optimization. On the official C-STS benchmark, PoLi-RL achieves a Spearman correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture. Furthermore, PoLi-RL also maintains SOTA status on the re-annotated C-STS dataset, confirming its robust generalization capabilities. As the first work to successfully apply RL to C-STS, our study introduces a powerful paradigm for aligning LLMs for complex, ranking-based conditional judgment tasks. Our code and checkpoints are available at https://github.com/ZBWpro/PoLi-RL.
@@ 1 Introduction
As a core research area in Computational Linguistics, Semantic Textual Similarity (STS) finds extensive applications across diverse scenarios, including topic modeling, dialogue systems, text summarization, and agent memory (Tang et al., 2025; Zhang et al., 2024b; 2025a). However, traditional STS tasks exhibit inherent ambiguity because similarity definitions are often susceptible to subjective bias. To address this limitation, the Conditional Semantic Textual Similarity (C-STS) task was developed (Deshpande et al., 2023). By incorporating an explicit natural language condition, C-STS enables more precise and objective similarity judgments, yet simultaneously imposes higher demands on a model’s reasoning capabilities. For instance, consider the following two text fragments: “A player is shooting from beyond the three-point line” and “A player is taking a free throw”. Under the condition “The activity of the player”, their similarity is high. However, under the condition “The player’s distance from the basket”, their similarity is low.
@@
Research on this nascent task has yielded three primary paradigms: Bi-encoder (Liu et al., 2025), Tri-encoder (Lin et al., 2024), and Cross-encoder (Li et al., 2024). The Cross-encoder architecture, which processes the text pair and the guiding condition simultaneously, is inherently compatible with modern generative pre-trained models. Despite this, the integration of C-STS with LLMs remains in its early stages. Current LLM applications are limited to two main approaches: direct inference via few-shot prompting, where even state-of-the-art models struggle to achieve satisfactory results (Deshpande et al., 2023); and utilizing them as feature extractors for generating text embeddings (Yamada & Zhang, 2025), which is an extension of the discriminative paradigm. To the best of our knowledge, no prior work has implemented an end-to-end LLM-based cross-encoder for the C-STS task, nor has any integrated it with advanced training techniques like reinforcement learning (RL), leaving a significant research gap.
@@
This paper aims to fill this gap. We argue that incorporating RL into an LLM-based cross-encoder paradigm is a natural fit. This is reflected in two key aspects: First, C-STS requires sophisticated, scenario-based reasoning. For example, in the basketball case described earlier, to correctly assess similarity under the “distance” condition, the model must transcend surface-level semantics to identify the underlying spatial relationship between ‘beyond the three-point line’ and ‘at the free throw line’, a process demanding strong abstraction and inference. RL, through its explicit reward signals, can more effectively guide the reasoning process of LLMs (Guo et al., 2025). Second, from an optimization standpoint, RL aligns closely with the task’s evaluation criteria. The Spearman correlation coefficient (Zar, 2005), a core evaluation metric of C-STS, is a non-differentiable measure of ranking quality. Traditional SFT methods can only indirectly and approximately optimize this objective through loss functions like Mean Squared Error (MSE) (Zhang & Li, 2024b). In contrast, RL allows for the direct optimization of ranking-based reward functions designed to correlate strongly with the final Spearman metric, aligning the training objective more closely with the evaluation criteria.
@@
However, a naive application of RL to this task presents significant challenges. As illustrated in Figure 1, our preliminary experiments indicate that directly applying a listwise ranking reward (e.g., Spearman’s correlation coefficient) across an entire batch of completions does not show any advantages compared to the few-shot method. This approach suffers from two fundamental problems. First, the ranking objective is too complex for a model that has not yet grasped the task’s fundamental scoring rules, often leading to training collapse. Second, a single reward computed across the entire batch is too coarse to provide precise credit assignment, as a few poor completions can unfairly penalize other good ones.
@@
Figure 1: Performance comparison of different strategies on the C-STS task. Directly applying a listwise ranking reward for RL does not significantly outperform the few-shot baseline. In contrast, both stages of our method (PoLi-RL) achieve substantial improvements, validating its effectiveness.
@@
To address these challenges, we propose PoLi-RL, a two-stage Point-to-List Reinforcement Learning framework. PoLi-RL features a two-stage curriculum to manage the complexity of the learning task. In the first stage, we use a simple pointwise reward to ground the model in the basic scoring rules of the task. Building on this foundation, the second stage introduces a richer, hybrid reward signal that combines a robust pointwise anchor with more nuanced pairwise and listwise ranking rewards, enabling the model to discern subtle semantic differences.
@@
Furthermore, to resolve the problem of coarse-grained reward signals that arises from ranking all completions in a batch together, we innovatively introduce a Parallel Slice Ranking Reward (PSRR) mechanism, which employs a two-level decomposition. First, for a batch of $N$ input samples, the model generates $G$ completions for each. We then form $G$ “parallel slices”, where the $i$-th slice is composed of the $i$-th completion from every sample. Second, and more importantly, within each slice, rather than assigning a single reward, we compute the rank difference for each individual completion against its ideal rank. This two-level decomposition allows each of the $N\times G$ completions to receive a unique and precise reward that reflects its quality, thereby enabling granular credit assignment and more effective optimization.
@@
  • To the best of our knowledge, this is the first work to propose an end-to-end, LLM-based cross-encoder for the C-STS task, and also represents the first application of reinforcement learning in this domain.
  • We design and implement PoLi-RL, a novel two-stage training curriculum that overcomes the optimization difficulties of direct rank-based learning by progressing from a basic pointwise objective to a comprehensive hybrid reward.
  • We propose the Parallel Slice Ranking Reward (PSRR) mechanism, which delivers precise and differentiated learning signals by computing ranking rewards within independent ‘parallel slices.’ This mechanism offers a generalizable strategy for other ranking tasks involving multiple candidate generations.
  • On the official C-STS benchmark, PoLi-RL achieves a Spearman’s correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture and surpassing strong proprietary models, including GPT-4o (44.23) and Deepseek-R1 (42.85). Furthermore, achieving SOTA on the re-annotated dataset and our qualitative analysis jointly confirm the model’s superior capability in nuanced conditional reasoning.
@@ 2 Methodology
This section details our proposed strategy. We begin by formulating the C-STS task within an end-to-end, LLM-based cross-encoder paradigm in subsection 2.1. Then, in subsection 2.2, we map this task onto the mathematical framework of Reinforcement Learning and specify its optimization objective. Finally, in subsection 2.3, we provide a comprehensive description of our PoLi-RL framework, including its two-stage design and the innovative PSRR mechanism.
@@ 2.1 Problem Formulation
The core objective of C-STS is to learn a scoring function that accurately reflects the semantic similarity between two text segments under a specific condition. Formally, each C-STS data sample is defined as a pair $(x,y)$, where the input $x=(t_{1},t_{2},c)$ consists of two text segments and a natural language condition, and the label $y\in[1,5]$ is the human-annotated similarity judgment on the Likert scale (Likert, 1932). Specifically, the label $y$ corresponds to a fine-grained set of semantic criteria. According to the C-STS annotation guidelines (Deshpande et al., 2023), the scores represent: (1) Completely dissimilar; (2) Thematically related but dissimilar; (3) Roughly equivalent, but with some important information differences; (4) Mostly equivalent, differing only in unimportant details; (5) Completely equivalent. This level of granularity demands that the model perform fine-grained reasoning beyond surface-level semantics, posing a significant challenge to its capabilities.
A distinctive feature of the C-STS dataset is its paired structure: samples are organized in adjacent pairs that share the same text segments $(t_{1},t_{2})$ but feature different conditions and maintain a strict ordinal relationship between their labels, i.e., $y_{\text{high}}\geq y_{\text{low}}$. This structure provides a solid foundation for our pairwise reward design, as detailed in subsection 2.3.
Our task is to train a scoring model $\pi_{\theta}$, parameterized by $\theta$. For each sample, the model takes a unified prompt $p=[\mathcal{I},\mathcal{E},x]$ (detailed in Appendix A.2) as input, consisting of the instruction $\mathcal{I}$, $K$ few-shot demonstrations $\mathcal{E}=\{(x_{k},y_{k})\}_{k=1}^{K}$, and the query input $x$, to generate an output sequence $o=\pi_{\theta}(p)$. From this sequence, we parse the final predicted score, $\tilde{y}=\text{Parse}(o)$. The overall training objective is to maximize the ranking consistency between the set of predicted scores $\{\tilde{y}_{i}\}_{i=1}^{N}$ and the ground-truth scores $\{y_{i}\}_{i=1}^{N}$ by optimizing the policy $\pi_{\theta}$, a process primarily quantified by Spearman’s correlation coefficient. Since this metric is rank-based and non-differentiable, RL emerges as a more promising optimization paradigm than traditional supervised fine-tuning.
@@ 2.2 Reinforcement Learning for C-STS
We formulate the C-STS task as a Markov Decision Process (MDP), defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where the agent is the LLM policy $\pi_{\theta}$. The generation process is formulated as a sequential decision-making process, where each step involves generating a single token. A state $s_{t}\in\mathcal{S}$ at timestep $t$ is the sequence of tokens generated so far, conditioned on the initial prompt, i.e., $s_{t}=(p,o_{<t})$. An action $a_{t}\in\mathcal{A}$ is the selection of the next token $o_{t}$ from the model’s vocabulary, governed by the policy $\pi_{\theta}(a_{t}|s_{t})$, which provides a probability distribution over all possible tokens. The transition function $\mathcal{T}$ is deterministic, where the next state $s_{t+1}$ is formed by appending the selected token $a_{t}$ to $s_{t}$. We employ a terminal reward setting, where a reward $\mathcal{R}_{T}=\mathcal{R}(x,y,o)$ is given only after the entire sequence $o$ has been generated. Finally, the discount factor $\gamma$ is set to 1 to ensure that the terminal reward is backpropagated without decay to all actions that contributed to the final output. Based on this framework, the objective is to find the optimal parameters $\theta^{*}$ that maximize the expected reward over the data distribution $\mathcal{D}$:
$\theta^{*}=\underset{\theta}{\operatorname{arg\,max}}\mathbb{E}_{(x,y)\sim\mathcal{D},o\sim\pi_{\theta}(p)}[R(x,y,o)]$ (1)
To optimize this objective, we employ Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), an extension of GRPO (Shao et al., 2024) that introduces several key techniques for effective RL. For each sample $x$, the policy model generates a set of $G$ completions $\{o_{i}\}_{i=1}^{G}$. A scalar reward $r_{i}=R(x,y,o_{i})$ is computed for each completion. The advantage $\hat{A}_{i}$ for each completion is then calculated by normalizing its reward against the statistics of the group rewards via Z-score normalization:
@@
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{(x,y)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|p)}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left(\frac{\pi_{\theta}(o_{i,t}|p,o_{i,<t})}{[\pi_{\theta}(o_{i,t}|p,o_{i,<t})]_{\text{nograd}}}\hat{A}_{i}\right)\right]$ (3)
where $\cdot]_{\text{nograd}}$ stops gradient flow through the denominator, ensuring only the numerator is updated.